# Super-Tonks-Girardeau quench of dipolar bosons in a one-dimensional optical lattice

### v1

@@ Abstract
A super-Tonks-Giradeau gas is a highly excited yet stable quantum state of strongly attractive bosons confined to one dimension. This state can be obtained by quenching the interparticle interactions from the ground state of a strongly repulsive Tonks-Girardeau gas to the strongly attractive regime. While the super-Tonks-Girardeau quench with contact interactions has been thoroughly studied, less is known about the stability of such a procedure when long-range interactions come into play. This is a particularly important question in light of recent advances in controlling ultracold atoms with dipole-dipole interactions. In this study, we thus simulate a super-Tonks-Girardeau quench on dipolar bosons in a one-dimensional optical lattice and investigate their dynamics for many different initial states and fillings. By calculating particle density, correlations, entropy measures, and natural occupations, we establish the regimes of stability as a function of dipolar interaction strength. For an initial unit-filled Mott state, stability is retained at weak dipolar interactions. For cluster states and doubly-filled Mott states, instead, dipolar interactions eventually lead to complete evaporation of the initial state and thermalization consistent with predictions from random matrix theory. Remarkably, though, dipolar interactions can be tuned to achieve longer-lived prethermal states before the eventual thermalization. Our study highlights the potential of long-range interactions to explore new mechanisms to steer and stabilize excited quantum states of matter.
@@ I Introduction
Rapid experimental advances in controlling ultracold atoms and their interactions are making it possible to realize and study enticing states of matter that encapsulate the exotic properties of quantum mechanics. These range from equilibrium phases such as supersolids and dipolar supersolids [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], to dynamical phenomena such as time crystals [15, 16, 17] and various Floquet phases [18, 19, 20, 21, 22, 23, 24]. Among these numerous recent advances, the super-Tonks-Girardeau (sTG) gas deserves special attention. The sTG is a quantum state of strongly attractive bosons confined to a one-dimensional (1D) geometry [25, 26, 27, 28, 29], and it is remarkable because it is stable despite being a highly excited state of matter. The sTG gas can be obtained from the standard Tonks-Girardeau (TG) gas [30, 31, 32, 33, 34, 35, 36, 37, 38] under sudden switching of contact interactions from the strongly repulsive to the strongly attractive regime. A possible explanation for the stability of the final state after the sudden quench relies on the sTG inheriting hardcore behavior from the TG gas [39] and the quenching process not causing any violent dynamics [40]. Besides providing a novel paradigmatic behavior, the realization of the sTG phase opens up the possibility of creating other quantum states of ultracold quantum gases with similar counter-intuitive properties [41, 42].
@@
In particular, in recent experiments with 1D arrays of bosons, the stability of the sTG phase has been revisited by adding a weak dipole-dipole interactions (DDIs) between the atoms. In the experiment of Kao et. al. [29], the robustness of the sTG phase against collapse has been studied for a bosonic one-dimensional quantum gas of dysprosium atoms. Unexpectedly, even though the DDIs break the integrability, they enhance the stability of the sTG phase in comparison to the short-range interaction case. This observation raises the prospect of employing weak dipolar forces to influence stability without changing the energy of the sTG phase significantly. This issue has been addressed recently by solving exactly a system of three trapped atoms with both contact and dipolar interactions [43], for which a distinct spectral response consisting of weaker hybridization with excited bound states is observed when repulsive DDIs are present. Another recent work [44] explores instead the unconventional mechanism of liquid formation in 1D optical lattice with strong on-site repulsion and weak long-range interactions. The underlying mechanism for liquid formation is established by a superexchange process. This discovery further prompted the investigation of the sTG quench in the corresponding geometry by means of extended Bose-Hubbard models, exploring the disruption of the states within a specific range of interactions [45].
@@
Our work aims to describe a similar sTG quench in an optical lattice but with richer initial setups spanning more diverse states and particle fillings. This allows us to unravel intriguing physics in the quench dynamics. We mainly focus on a) clean Mott states and cluster states with unit filling and b) fermionized Mott states and cluster states with double filling. For each case, the strong short-range on-site repulsion is supplemented by four (for unit filling) and three (for double filling) specific cases of long-range dipolar interactions which correspond to different scenarios in the dynamical process. We present numerically exact many-body dynamics by solving the time-dependent many-boson Schrödinger equation utilizing the MultiConfigurational Time-Dependent Hartree method for bosons (MCTDHB) [46, 47, 48, 49, 50, 51, 52], implemented in the MCTDH-X software [53, 54]. We extract key measures such as one- and two-body density dynamics, natural occupations, many-body information entropy and Glauber correlation functions. Our many-body dynamics highlights microscopic mechanisms not explored in earlier studies and provide understanding of the pathway from TG to sTG transition.
Besides highlighting new pathways to stable excited states of matter, studies on DDIs in STG states – such as the experiment by Kao et. al [29] – also give rise to strongly correlated prethermal states. In fact, prethermalization has emerged over the past decade as a ubiquitous phenomenon in the dynamics of ultracold quantum gases in 1D geometries [55, 56, 57, 58, 59, 60]. Prethermalization happens when a many-body quantum system exhibits non-monotonic non-equilibrium dynamics: before reaching the true relaxed state (e.g. a thermal state), the system first relaxes to a short lived prethermal state which exhibits well separated intermediate time scales manifested by different values of certain observables, first among all in entropic quantities. This is contrast to the monotonic way of relaxation exhibiting a single time scale. While investigating the relaxation process in our STG quench dynamics, we encounter a highly nontrivial time evolution that leads to prethermalization. In particular, we show how to control the long-range interactions to create a long-lived prethermal states or short-lived quasi-prethermal states.
Our overall observations are as follows: While the TG gas without dipolar perturbation is stable under the sTG quench, all weakly bound and scattering Mott states and all cluster states collapse and lose all correlations – a process termed evaporation. The evaporation occurs at different time scales depending on the strength of the DDIs and we can trace clear many-body features before it actualizes. Furthermore, the fermionized Mott state exhibits even more intriguing expansion dynamics before evaporation. In terms of entropy dynamics, perfect thermalization is observed both for the clean Mott states and deep self-bound cluster states. However, for very weak attractive long-range interactions, some quasi-stationary states at intermediate time scales emerge. This establishes how long-range interactions can be taken as a control parameter to create prethermal or quasi-prethermal states with longer lifetimes.
@@ II System and protocol
Its one-body part is $\hat{h}(x)=\hat{T}(x)+\hat{V}_{OL}(x)$, where $\hat{T}(x)=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}}{\partial x^{2}}$ is the kinetic energy operator and $\hat{V_{OL}}(x)=V_{0}\sin^{2}(k_{0}x)$ is the external 1D lattice of depth $V_{0}$ and wave vector $k_{0}$. Unless otherwise stated, we will keep $V_{0}$ fixed at 20 $E_{r}$, where $E_{r}=\frac{\hbar k_{0}}{2m}$ is the recoil energy of the lattice. This depth is necessary to reach the Mott regime in the state preparation. We restrict the geometry to encompass $S$ sites in the optical lattice (defined as the spatial extent between two maxima in the sinusoidal function) adding hard-wall boundary conditions at each end. The two-body interactions $\hat{W}(x_{i}-x_{j})$ contain both short-range $W_{C}(x_{i},x_{j})=g_{0}\delta(x_{i}-x_{j})$ and long-range dipolar interactions $W_{D}(x_{i},x_{j})=\frac{g_{d}}{|x_{i}-x_{j}^{\prime}|^{3}+\alpha}$. Here, $g_{0}$ is the strength of the local interactions, which is strongly positive before the quench and suddenly becomes negative after the quench, $g_{d}$ controls the strength of the long-range interactions and it is negative all throughout our calculations, and $\alpha$ denotes a renormalization factor used to avoid nonphysical singularities at $x_{i}=x_{j}$. Without long-range interactions, the ground state of the system is either in the TG phase (with one particle per site) or the fermionized TG phase (with more than one particle per site). With strong negative DDIs, the ground state is a cluster state at one site in the optical lattice and possesses distinguished many-body features. Weakly negative DDIs interpolate between these two cases. We remark that the Hamiltonian $\hat{H}$ can be written in dimensionless units obtained by dividing the dimensionful Hamiltonian by $\frac{\hbar^{2}}{mL^{2}}$, with $L$ an arbitrary length scale, which for the sake of our calculations we will set to be the period of the optical lattice.
@@
With this Hamiltonian, we will study two different setups. First, we consider a unit filling scenario of $N=5$ particles in $S=5$ sites. Then, we explore a more correlated double filling setup, with $N=6$ particles in $S=3$ sites. The protocol in both cases is identical. We first obtain the ground state for the Hamiltonian with strong contact repulsion ($g_{0}^{t<0}\gg 0$). Then, at time $t=0$, we propagate this state by quenching it to the strongly attractive regime of equal magnitude ($g_{0}^{t\geq 0}\ll 0$, $|g_{0}^{t\geq 0}|=g_{0}^{t<0}$). In the experiments, this is achieved by exploiting a confinement-induced resonance (CIR), where particle interactions are enhanced in a confined space such as a trap or potential well [33, 34, 61, 62, 63, 64, 65, 66]. By means of Feshbach resonance techniques, it is possible to control the $s$-wave contact interaction and effectively control the long-range interaction. We study the stability of the sTG state during the time evolution for different values of the attractive dipolar interactions $g_{d}$. Our additional motivation is to find the time scale of relaxation and possibly confirm the existence of any prethermal states.
@@ III Methods
To investigate the dynamics of dipolar bosons under the sudden quench, we employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles (MCTDHB) [46, 47, 48, 49] implemented by the MCTDH-X software [49, 50, 51, 53, 52, 54]. MCTDH-X solves the many-body Schrödinger equation by recasting the many-body wave function as an adaptive superposition of $M$ time-dependent single-particle wave functions, called orbitals. Both the coefficients and the basis functions in this superposition are optimized in time to yield either ground-state information (via imaginary time propagation) or full-time dynamics (via real-time propagation). We refer to the Appendix  A for further details on this method. In our calculations, we employ $M=10$ orbitals for the unit filling case of $N=5$ particles in $S=5$ sites, and $M=12$ orbitals for the double filling case of $N=6$ particles in $S=3$ sites. Our choice of number of orbitals thus guarantees an average of at least two orbitals per site in each case and the required convergence is assured.
To extract information from the system, we calculate several observables. To probe the spatial distribution of the bosons, we calculate the one-body density from the many-body state $\left|\Psi(t)\right>$ as
@@
To study statistical relaxation and thermalization in the quench dynamics, the information theoretical measure known as Shannon information entropy is the ideal quantity [67, 68]. We define here an equivalent measure of information entropy called occupation entropy. This is based on the assumption that in MCTDHB with time evolution the many-body state can exhibit dynamical fragmentation, i.e. several natural orbitals exhibit significant population. We thus define the occupation information entropy as
$S_{\vec{n}}(t)=-\sum_{\vec{n},i}\frac{n_{i}(t)}{N}\ln\left(\frac{n_{i}(t)}{N}\right),$ (6)
where $n_{i}(t)$ is the occupation of the $i$-th element of the many-body configuration $\vec{n}$ (see appendix A for details). In the Gross-Pitaevskii single orbital mean-field theory, as the reduced density matrix has a single eigenvalue, $S_{\vec{n}}$ is zero. For multiconfigurational states, instead, when several configurations contribute to the many-body state, $S_{\vec{n}}\neq 0$. As more and more orbitals start to populate the many-body state, $S_{\vec{n}}(t)$ gradually increases and finally saturates to a value determined by the maximally fragmented many body state.
@@ IV Results for unit filling
We begin our analysis by presenting the results for the unit filling case of $N$=5 bosons in $S=5$ sites. In this section, unless otherwise stated, the magnitude of the contact interactions is fixed at $|g_{0}|\approx 81E_{r}$. We have probed a fine array of dipolar interactions in the range $g_{d}\in[-3.6E_{r},0E_{r}]$, but we will only show representative results to be concise.
IV.1 Evaporation mechanism
@@
Figure 1: Density dynamics post quench for $N=5$ bosons and $M=10$ orbitals in $S=5$ sites (unit filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.45E_{r}$, (c) $g_{d}=-0.82E_{r}$, (d) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$. The dotted lines indicate the system boundaries.
The density dynamics is plotted in Fig. 1(a)-(d) for four representative values of $g_{d}$ showcasing the appearance of different dynamical behaviors. Fig. 1(a) demonstrates the TG to sTG quench for $g_{d}=0$. In this scenario, the pre-quench state is a fully fragmented clean Mott state in the TG limit (see appendix C for an overview of the initial states), which has positive energy and remains highly stable for very long time after the quench. Upon turning on the dipolar attractions, for $g_{d}=-0.45E_{r}$, the bosons initially retain their Mott structure. However, the post-quench stability is reduced and the particles eventually evaporate as shown in Fig. 1(b). When the long-range attractions are increased further, e.g. to $g_{d}=-0.82E_{r}$, the bosons progressively deviate from the Mott state and form clusters pinned around a subset of the total number of lattice sites. The system still has positive energy, but due to the effect of long-range interactions all the five bosons are confined within a subset of five lattice sites. Upon quenching, the state first demonstrates very intriguing many-body expansion dynamics and subsequently evaporates as shown in Fig. 1(c). We remark that the time scale for this process is extremely short (evaporation begins already around 10 $\mu$s) compared to the time scale for the emergence of instability in the clean Mott state. For even stronger attractions, which correspond to the self bound Mott phase with negative energy, the bosons form a cluster state in the central lattice and upon quenching the evaporation is extremely fast – in the order of 5 $\mu s$. This is illustrated in Fig. 1(d) for $g_{d}=-2.02E_{r}$. In both Fig. 1(c) and (d), the diverging nature of long-range interactions is clear: the density is propelled externally, hits the walls of the lattice, bounces back, and forms a distorted interference pattern.
@@
Figure 2: Dynamics of the reduced one-body density $\rho^{(1)}(x,x^{\prime})$ post quench for $N=5$ bosons and $M=10$ orbitals in $S=5$ sites (unit filling) with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.45E_{r}$ (g)-(i) $g_{d}=-0.82E_{r}$, (j)-(l) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.

### v2

# Interaction quench of dipolar bosons in a one-dimensional optical lattice

@@ Abstract
A Tonks-Girardeau (TG) gas is a highly correlated quantum state of strongly interacting bosons confined to one dimension, where repulsive interactions make the particles behave like impenetrable fermions. By suddenly tuning these interactions to the attractive regime, it is possible to realize a super-Tonks-Girardeau (sTG) gas – a highly excited, metastable state of strongly attractive bosons with unique stability properties. Inspired by the sTG quench scenario, we investigate a similar setup but with the inclusion of long-range dipolar interactions, which modify the system away from the TG Mott insulating limit. We simulate an interaction quench on dipolar bosons initially prepared in various states and fillings, using real-space densities, orbital occupations, Glauber correlation functions, and autocorrelation functions to probe post-quench stability. Our results reveal that stability is maintained only at very weak dipolar interaction strengths when starting from a unit-filled TG Mott state. In contrast, all cluster states – whether unit-filled or doubly-filled – eventually collapse under attractive interactions. This collapse is not always visible in the density profile but becomes apparent in the autocorrelation function, indicating complex many-body restructuring of the quantum state. Our findings underscore the potential of dipolar interactions to drive novel quantum dynamics and highlight the delicate balance required to stabilize excited states in long-range interacting systems.
@@ I Introduction
Rapid experimental advances in controlling ultracold atoms and their interactions are making it possible to realize and study enticing states of matter that encapsulate the exotic properties of quantum mechanics. These range from equilibrium phases such as supersolids and dipolar supersolids [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], to dynamical phenomena such as time crystals [15, 16, 17] and various Floquet phases [18, 19, 20, 21, 22, 23, 24]. Among these numerous recent advances, the super-Tonks-Girardeau (sTG) gas deserves special attention. The sTG is a quantum state of strongly attractive bosons confined to a one-dimensional (1D) geometry [25, 26, 27, 28, 29], and it is remarkable because it is metastable despite being a highly excited state of matter [30, 31]. The sTG gas can be obtained from the standard Tonks-Girardeau (TG) gas [32, 33, 34, 35, 36, 37, 38, 39, 40, 41] under sudden switching of contact interactions from the strongly repulsive to the strongly attractive regime. The sTG state resembles a highly excited gas-like state and, in principle, it should have an infinite life time as it is an eigenstate of the Hamiltonian. Experimentally, it is in fact demonstrated that for short-range interactions the sTG state is a long-lived excitation. Besides providing a novel paradigmatic behavior, the realization of the sTG phase opens up the possibility of creating other quantum states of ultracold quantum gases with similar counter-intuitive properties [42, 43, 44].
@@
In particular, in recent experiments with 1D arrays of bosons, the stability of the sTG phase has been revisited by adding a weak dipole-dipole interactions (DDIs) between the atoms. In the experiment of Kao et. al. [29], the robustness of the sTG phase against collapse has been studied for a bosonic one-dimensional quantum gas of dysprosium atoms. Unexpectedly, even though the DDIs break the integrability, they enhance the stability of the sTG phase in comparison to the short-range interaction case. This observation raises the prospect of employing weak dipolar forces to influence stability without changing the energy of the sTG phase significantly. This issue has been addressed recently by solving exactly a system of three trapped atoms with both contact and dipolar interactions [45], for which a distinct spectral response consisting of weaker hybridization with excited bound states is observed when repulsive DDIs are present. Another recent work [46] explores instead the unconventional mechanism of liquid formation in 1D optical lattice with strong on-site repulsion and weak long-range interactions. The underlying mechanism for liquid formation is established by a superexchange process. This discovery further prompted the investigation of the sTG quench in the corresponding geometry by means of extended Bose-Hubbard models, exploring the disruption of the states within a specific range of interactions [47]. In this setup, strong local interactions are supplemented by weaker and attractive nonlocal interactions which lead to i) a gas phase, ii) a liquid phase and a iii) a self bound Mott phase. In the post quench dynamics, the liquid droplets expand and eventually evaporate.
Our work aims to describe a similar interaction quench in an experimentally realistic continuum realization with bosons trapped in an optical lattice. We study richer initial setups, spanning more diverse states and particle fillings. This allows us to unravel unexplored many-body mechanisms in the dynamics. We characterize the post quench metastable states with specific calculations of their life times and measures of correlation and many-body coherence. We also elaborate on how the stability of the state can be controlled by the additional weak dipolar interactions.
@@
Our initial conditions span clean Mott states, cluster states, and single-site localized quantum droplets with a) unit filling and b) double filling. For each case, the short-range on-site repulsion is supplemented by long-range dipolar interactions which correspond to different initial states and dynamical scenarios. We present many-body dynamics results by solving the time-dependent many-boson Schrödinger equation utilizing the MultiConfigurational Time-Dependent Hartree method for bosons (MCTDHB) [48, 49, 50, 51, 52, 53, 54], implemented in the MCTDH-X software [55, 56]. We extract key measures such as one- and two-body density dynamics, natural occupations, one- and two-body Glauber correlation functions, and autocorrelation functions.
Our overall observations are as follows: The TG gas without dipolar perturbation and with very weak dipolar attractions – which correspond to scattering states – remains stable for a long time under the interaction quench, as expected. However, in contrast to the findings in Bose-Hubbard models, all weakly bound Mott states and cluster states eventually deviate from the initial state and even collapse completely. This is particularly true for fragmented quantum droplets where all the particles are concentrated in the same site. The one-body and two-body correlations reveal the details of this destabilization, with progressive loss of coherence, loss of correlation holes, and increase of correlation squeezing. Our findings further reinforce the uniqueness of sTG states in their stability and contextualize the effect of long-range interactions in a more realistic continuum description.
@@ II System and protocol
Its one-body part is $\hat{h}(x)=\hat{T}(x)+\hat{V}_{OL}(x)$, where $\hat{T}(x)=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}}{\partial x^{2}}$ is the kinetic energy operator and $\hat{V_{OL}}(x)=V_{0}\sin^{2}(k_{0}x)$ is the external 1D lattice of depth $V_{0}$ and wave vector $k_{0}$. We remark that the Hamiltonian $\hat{H}$ can be written in dimensionless units obtained by dividing the dimensionful Hamiltonian by $\frac{\hbar^{2}}{mL^{2}}$, with $L$ an arbitrary length scale, which for the sake of our calculations we will set to be the period of the optical lattice. Unless otherwise stated, we will keep $V_{0}$ fixed at 20 $E_{r}$, where $E_{r}=\frac{\hbar^{2}k_{0}^{2}}{2m}$ is the recoil energy of the lattice. This depth is necessary to reach the Mott regime in the state preparation. We restrict the geometry to encompass $S$ sites in the optical lattice (defined as the spatial extent between two maxima in the sinusoidal function) adding hard-wall boundary conditions at each end. The two-body interactions $\hat{W}(x_{i}-x_{j})$ contain both short-range $W_{C}(x_{i},x_{j})=g_{0}\delta(x_{i}-x_{j})$ and long-range dipolar interactions $W_{D}(x_{i},x_{j})=\frac{g_{d}}{|x_{i}-x_{j}^{\prime}|^{3}+\alpha}$. Here, $g_{0}$ is the strength of the local interactions, which is strongly positive before the quench and suddenly becomes negative after the quench, $g_{d}$ controls the strength of the long-range interactions and it is negative all throughout our calculations, and $\alpha$ denotes a renormalization factor used to avoid nonphysical singularities at $x_{i}=x_{j}$. Without long-range interactions, the ground state of the system is either in the TG Mott phase (with one particle per site) or in a multiple-filled Mott phase (with more than one particles per site). With strong negative DDIs, the ground state is a self-bound quantum droplet state at the central site in the optical lattice and possesses distinguished many-body features. Weakly negative DDIs interpolate between these two cases, creating variations of cluster or droplet states where the central site has strong (potentially multiple) occupation but the neighboring sites see a rapid decrease in population.
@@
With the Hamiltonian above, we will study two different setups. First, we consider a unit filling scenario of $N=5$ particles in $S=5$ sites. Then, we explore a more correlated double filling setup, with $N=6$ particles in $S=3$ sites. The protocol in both cases is identical. We first obtain the ground state for the Hamiltonian with strong contact repulsion ($g_{0}^{t<0}\gg 0$). Then, at time $t=0$, we propagate this state by quenching it to the strongly attractive regime of equal magnitude ($g_{0}^{t\geq 0}\ll 0$, $|g_{0}^{t\geq 0}|=g_{0}^{t<0}$). In the experiments, this is achieved by exploiting a confinement-induced resonance (CIR), where particle interactions are enhanced in a confined space such as a trap or potential well [35, 36, 57, 58, 59, 60, 61, 62]. By means of Feshbach resonance techniques, it is possible to control the $s$-wave contact interaction and indirectly also control the long-range interactions. We study the stability of the different initial states – ranging from sTG to droplet – after the sudden quench for different values of the attractive dipolar interactions $g_{d}$.
@@ III Methods
To investigate the dynamics of dipolar bosons under the sudden quench, we employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles (MCTDHB) [48, 49, 50, 51] implemented by the MCTDH-X software [51, 52, 53, 55, 54, 56]. MCTDH-X solves the many-body Schrödinger equation by recasting the many-body wave function as an adaptive superposition of $M$ time-dependent permanents constructed from single-particle wave functions, called orbitals. Both the coefficients and the basis functions in this superposition are optimized in time to yield either ground-state information (via imaginary time propagation) or full-time dynamics (via real-time propagation). We refer to the Appendix  A for further details on this method.
Utilizing several orbitals in the many-body ansatz allows us to capture fragmented many-body states, i.e. states that cannot be described by mean-field approximations, like sTG states. In general, depending on the quench process and the quenched parameter of the initial Hamiltonian, fragmentation may increase (or less typically decrease) with time – a phenomenon usually called dynamical fragmentation. This prompts the question of orbital convergence for a fragmented state, i.e. how many orbitals are required to capture the dynamics correctly. In MCTDHB, one method to ensure convergence is by repeating the computation with a higher number of orbitals until the dynamical measures converge and the occupation in the last orbital becomes insignificant. For very strong quenches, obtaining the exact, converged dynamics at long times might require a computationally unfeasible number of orbitals; therefore, only short and intermediate time dynamics can be obtained reliably. For moderately strong quenches like the ones we are probing in this work, we are able to construct a converged dynamics during the entire time evolution by employing up to $M=15$ orbitals. Details on the calculations that confirm the orbital convergence of our results are offered in appendix C.
To extract information from the system, we calculate several observables from the many-body state $\left|\Psi(t)\right>$. To probe the spatial distribution of the bosons, we calculate the one-body density as
@@
In appendix E, we also showcase results pertaining to normalized one-body and two-body densities – quantities normally referred to as one-body and two-body Glauber correlation functions $g^{(1)}(x,x^{\prime})=\frac{\rho^{(1)}(x,x^{\prime})}{N\sqrt{\rho(x)\rho(x^{\prime})}}$ and $g^{(2)}(x,x^{\prime})=\frac{\rho^{(2)}(x,x^{\prime})}{N^{2}\rho(x)\rho(x^{\prime})}$.
From the reduced one-body density matrix, it is possible to obtain information regarding the orbital occupation throughout the time evolution, i.e. how much of the provided Hilbert space is dynamically occupied. This can be expressed via the natural orbitals $\phi_{i}^{(\mathrm{NO})}$ and the orbital occupations $\rho_{i}$, which are the eigenfunctions and eigenvalues of $\rho^{(1)}(x,x^{\prime})$, i.e.
$\rho^{(1)}(\mathbf{x},\mathbf{x}^{\prime})=\sum_{i}\rho_{i}\phi^{(\mathrm{NO}),*}_{i}(\mathbf{x}^{\prime})\phi^{(\mathrm{NO})}_{i}(\mathbf{x}).$ (6)
The dynamics of the orbital occupations offer another instrument to quantify the stability (or lack thereof) of the initial state after the quantum quench.
Finally, to study the eventual fate of all the different initial states – from the TG Mott states to the single-site clusters – and verify their stability under the sTG quench, we also calculate the autocorrelation function
$\mathcal{F}(t)\equiv|\left<\Psi(0)\middle|\Psi(t)\right>|^{2}.$ (7)
This quantity offers an immediate, quantitative measure of the post-quench stability at the many-body level. If it remains close to 1 for long times after the quench, we can consider the initial state stable with respect to the sudden quench. A decay away from this value indicates instead a loss of coherence and the eventual destabilization of the initial state.
@@ IV Results for unit filling
We begin our analysis by presenting the results for the unit filling case of $N=5$ bosons in $S=5$ sites. In this section, unless otherwise stated, the magnitude of the contact interactions is fixed at $|g_{0}|\approx 2E_{r}$. We have probed a fine array of dipolar interactions in the range $g_{d}\in[-0.4E_{r},0E_{r}]$, but we will only show representative results to be concise.
IV.1 Density
@@
Figure 1: Density dynamics post quench for $N=5$ bosons and $M=15$ orbitals in $S=5$ sites (unit filling) with (a) $g_{d}=-0.16E_{r}$, (b) $g_{d}=-0.18E_{r}$, (c) $g_{d}=-0.19E_{r}$, (d) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$. The dotted lines indicate the system boundaries.
The density dynamics is plotted in Fig. 1(a)-(d) for four representative values of $g_{d}$ showcasing the appearance of different dynamical behaviors. All computations shown in the figures have been performed with $M=15$ orbitals. We have systematically checked convergence in appendix C from $M=12$ up to $M=15$ orbitals both for the initial states and for the time evolution, and determined that the occupation of the last few orbitals is negligible. This confirms that our results are converged in the number of orbitals throughout the dynamics.
The initial state retains a fully fragmented, clean TG Mott state configuration for all $g_{d}$ values down to around $g_{d}=-0.16E_{r}$. This state has positive energy and remains highly stable for very long time after the quench. Fig. 1(a) shows the density dynamics for the TG to sTG quench in this case. From a density perspective, we can appreciate the robustness of this state, which appears unaffected by the interaction quench.
At around $gd=-0.18E_{r}$, the initial density profile begins to morph into a cluster state, which can have non zero population on all five sites [Fig. 1(b), $gd=-0.18E_{r}$] or only on the center-most three [Fig. 1(c), $gd=-0.19E_{r}$]. The system still has positive energy, but – due to the effect of long-range interactions – the five bosons are not uniformly localized. Upon quenching, these cluster states overall retain their initial shape, but oscillations in the density of each peak and a slow but progressive accumulation of the density towards the center can be observed. This already signals a destabilization with respect to the initial configuration, as opposed to the pure TG Mott state. We remark that the parameter space for the observation of these intermediate cluster states is very narrow, in the order of 0.02 $E_{r}$. This is reminiscent of the tiny sliver in parameter space where a quantum liquid phase can be observed in extended Bose-Hubbard models [47].
For even stronger attractions $g_{d}<-0.2E_{r}$, the bosons form a self-bound droplet with negative energy in the central lattice. Upon quenching, the collapse of the initial state is much more pronounced. In the order of $\approx 70$ $\mu$s, the initial droplet is compressed to a much narrow range due to the strong attractions. This is illustrated in Fig. 1(d) for $g_{d}=-0.4E_{r}$.
@@
Figure 2: Dynamics of the reduced one-body density $\rho^{(1)}(x,x^{\prime})$ post quench for $N=5$ bosons and $M=15$ orbitals in $S=5$ sites (unit filling) with (a)-(c) $g_{d}=-0.16E_{r}$, (d)-(f) $g_{d}=-0.18E_{r}$, (g)-(i) $g_{d}=-0.19E_{r}$, (j)-(l) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.